jessica and her sister have ages that total 55. jessica's sister is one more than twice jessica age. write a system of equations to represent this scenario
Answer by FrankM(1040)

J+S=55
S=2J+1
(you only asked for equations, but here is the rest)
J+2J+1=55
3J+1=55
3J=54
J=18
S=37